OpenStudy (anonymous):

I have 2 homework questions that have really been giving me trouble.. if ANYONE can help with either I'd be so extremely grateful. 1) Find the global extrema of f(x)= (1/3)x^3 - (1/2)x^2 + 6 on the interval [-3,4] 2) A farmer encloses a rectangular area so that area=200ft^2. The farmer will enclose three sides as there is already a fence on the fourth side. What dimensions should the rectangle be to use the least amount of fence? How much fence will be used?

OpenStudy (anonymous):

Are you familiar with derivatives? You're going to need them for these two

OpenStudy (anonymous):

Doulikepiecauseidont: yes and no.....this is my first calculus class and the professor is horrible. I know how to do the derivatives but in these problems am unsure of where to go

OpenStudy (anonymous):

first of all you need to study derivatives because you don't expect us to answer your question because you won't understand a word we will say. It's easy once you learn it. Believe me. I've been there. and Done that.

OpenStudy (anonymous):

Twinyang, if you would have read my previous comment you would see that I am familiar with derivatives. However, the process of solving the (above) question is new to me (I.e. I need some direction). Further, seeing as this is a public mathematics "help forum" I came for just that...help. If I understood all the aforementioned words^ I wouldn't be here in the first place. Thank you for the helpful advice.

ganeshie8 (ganeshie8):

Notice that the given funciton is a polynomial, So the max/mins can occur only at two places : 1) when f'(x) = 0 2) at the given boundary : -3, 4

ganeshie8 (ganeshie8):

find the derivative of f(x), set it equal to 0 and solve x

ganeshie8 (ganeshie8):

\(\large f(x)= (1/3)x^3 - (1/2)x^2 + 6 \) \(\large f'(x)=? \)

OpenStudy (anonymous):

f'(x) = x^2 - x ?

ganeshie8 (ganeshie8):

yes ! set it equal to 0 and solve x

OpenStudy (anonymous):

So, I did that and I factored out an x, --> looking something like x(x-1) = 0 and I got x= 1 But I'm not sure if that is correct and/or where to go from there

ganeshie8 (ganeshie8):

x(x-1) = 0 gives two solutions : x = 0, 1 right ?

OpenStudy (anonymous):

^ i.e where the interval comes in to play.... yes, two solutions! I'm sorry

ganeshie8 (ganeshie8):

So the max/mins can only occur at below `x` values : 0, 1, -3, 4

ganeshie8 (ganeshie8):

evaluate the given function at these x values, then pick the MAX value for global max, and pick the MIN value for global min.

OpenStudy (anonymous):

Do I plug those in to the original f(x)? or f'(x)?

ganeshie8 (ganeshie8):

\(\large f(x)= (1/3)x^3 - (1/2)x^2 + 6 \) \(\large f(0) = ? \) \(\large f(1) = ? \) \(\large f(-3) = ? \) \(\large f(4) = ? \)

OpenStudy (anonymous):

f(0) = 6 f(1) = ~5.83 f(-3) = -7.5 f(4) = 19.33 So if these are correct then my global max would be at f(4) and the min at f(-3) ?

ganeshie8 (ganeshie8):

Exactly !

OpenStudy (anonymous):

YES! That is so exciting. Thank you so, so much. I've really been struggling with the end of this assignment!

ganeshie8 (ganeshie8):

glad to see you figured it out :) yw !

ganeshie8 (ganeshie8):

Oh there is a second problem... lets see..

OpenStudy (anonymous):

This is the problem:

ganeshie8 (ganeshie8):